# EconWebArena: Benchmarking Autonomous Agents on Economic Tasks in Realistic Web Environments

### v1

@@ Abstract
Zefang Liuthanks: These authors contributed equally to this work. Georgia Institute of Technology Atlanta, GA 30332, USA liuzefang@gatech.edu                        Yinzhu Quan Georgia Institute of Technology Atlanta, GA 30332, USA yquan9@gatech.edu
@@ 1 Introduction
Direct interaction with authoritative online sources is critical for high-fidelity economic data acquisition, but it poses unique challenges for autonomous agents. These agents must operate in live web environments, reason over both structured and visual content, and execute multi-step procedures to obtain specific, verifiable information. However, existing benchmarks rarely reflect these demands. Most web agent benchmarks (Zhou et al., 2023; Drouin et al., 2024; Yoran et al., 2024) focus on general-purpose tasks such as shopping, email handling, or navigating productivity tools. These emphasize routine interactions but overlook the structured reasoning, domain expertise, and precision needed in economic contexts.
@@ 2.1 Web-Based Benchmarks
Web-based agent benchmarks have primarily focused on general-purpose tasks such as navigation, data entry, and e-commerce, with foundational platforms like WebArena (Zhou et al., 2023) and BrowserGym (Chezelles et al., 2024) enabling large-scale or modular evaluation. Specialized benchmarks include WebShop (Yao et al., 2022) for online shopping and WorkArena (Drouin et al., 2024; Boisvert et al., 2024) for productivity tasks. More recent efforts such as AssistantBench (Yoran et al., 2024), Mind2Web (Deng et al., 2023), WebCanvas (Pan et al., ), RealWebAssist (Ye et al., 2025), and WebLINX (Lù et al., 2024) aim to test broader agent capabilities in cross-task generalization, long-horizon reasoning, and interactive assistance. To support multimodal grounding, benchmarks like VisualWebArena (Koh et al., 2024), VideoWebArena (Jang et al., 2024), and MMINA (Zhang et al., 2024c) incorporate image and video inputs, while VisualAgentBench (Liu et al., 2024c) and VisualWebBench (Liu et al., 2024b) assess agents’ ability to align language with UI elements. However, none of these settings focus on the specialized challenges of economic data retrieval and domain-specific reasoning.
@@ 3.3 Observation and Action Spaces
EconWebArena extends the BrowserGym [^3] https://github.com/ServiceNow/BrowserGym Chezelles et al. (2024) framework to support economic tasks in realistic web environments. Each task provides the agent with structured observations, including the accessibility tree (AXTree), webpage screenshots, and contextual metadata such as focused elements, error logs, and action history. These inputs enable both language-based and multimodal reasoning. The action space includes fine-grained browser control primitives at both the element and coordinate levels, supporting operations like mouse movement, text input, tab switching, and form submission. This setup allows agents to perform the low-level interactions required for complex economic data retrieval. Further details are provided in Section 4.1.
@@ 4.2.3 RQ3: What directions can improve LLM agent performance on these tasks?
Our error analysis suggests several directions to improve LLM agents on EconWebArena tasks. First, robust visual grounding and multimodal reasoning are essential for interpreting economic charts, tables, and diagrams, which frequently appear in official data portals. Structured input representations, such as combining AXTree with screenshots and coordinates, can better support perception and interaction. Planning and self-criticism prompting also improve performance in multi-step navigation tasks. In addition, hybrid methods that integrate GUI-based exploration with structured API calls, when available, can enhance both reliability and efficiency. Finally, domain-adaptive tuning and retrieval-augmented prompting using realistic economic examples may further strengthen agent accuracy in complex cases.
@@ B.1 More Experimental Settings
In this section, we provide additional configuration details for all experiments conducted on EconWebArena. Table 7 summarizes the prompt-level settings used by agents, adapted from the AgentLab (Chezelles et al., 2024) interface. Unless otherwise specified, these settings were fixed across all model evaluations. We evaluate a comprehensive set of LLMs, including GPT-4o666gpt-4o-2024-08-06: https://platform.openai.com/docs/models/gpt-4o (Achiam et al., 2023; OpenAI, 2024b), GPT-4o mini777gpt-4o-mini-2024-07-18: https://platform.openai.com/docs/models/gpt-4o-mini (OpenAI, 2024a), GPT-4.1888gpt-4.1-2025-04-14: https://platform.openai.com/docs/models/gpt-4.1 (OpenAI, 2025a), GPT-4.1 mini999gpt-4.1-mini-2025-04-14: https://platform.openai.com/docs/models/gpt-4.1-mini (OpenAI, 2025a), o4-mini101010o4-mini-2025-04-16: https://platform.openai.com/docs/models/o4-mini (OpenAI, 2025b), Claude Sonnet 4111111claude-sonnet-4-20250514: https://openrouter.ai/anthropic/claude-sonnet-4 (Anthropic, 2025b), Gemini 2.5 Flash121212gemini-2.5-flash-preview-05-20: https://openrouter.ai/google/gemini-2.5-flash-preview-05-20 (Gemini Team, 2023; Google DeepMind, 2025), and Llama 4 Maverick131313llama-4-maverick-17b-128e-instruct-fp8: https://openrouter.ai/meta-llama/llama-4-maverick (Grattafiori et al., 2024; Meta AI, 2025).

### v2

@@ Abstract
Zefang Liuthanks: These authors contributed equally to this work. Capital One San Jose, CA, USA zefang.liu@capitalone.com          Yinzhu Quan Georgia Institute of Technology Atlanta, GA, USA yquan9@gatech.edu
@@ 1 Introduction
Direct interaction with authoritative online sources is critical for high-fidelity economic data acquisition, but it poses unique challenges for autonomous agents. These agents must operate in live web environments, reason over both structured and visual content, and execute multi-step procedures to obtain specific, verifiable information. In practice, such tasks mirror typical workflows in applied economics, for example retrieving official Consumer Price Index (CPI) releases for inflation analysis, collecting central bank interest rate data for policy evaluation, or accessing trade and labor statistics for empirical research. However, existing benchmarks rarely reflect these demands. Most web agent benchmarks (Zhou et al., 2023; Drouin et al., 2024; Yoran et al., 2024) focus on general-purpose tasks such as shopping, email handling, or navigating productivity tools. These emphasize routine interactions but overlook the structured reasoning, domain expertise, and precision needed in economic contexts.
@@ 2.1 Web-Based Benchmarks
Web-based agent benchmarks have primarily focused on general-purpose tasks such as navigation, data entry, and e-commerce, with foundational platforms like WebArena (Zhou et al., 2023) and BrowserGym (Chezelles et al., 2024) enabling large-scale or modular evaluation. Specialized benchmarks include WebShop (Yao et al., 2022) for online shopping and WorkArena (Drouin et al., 2024; Boisvert et al., 2024) for productivity tasks. More recent efforts such as AssistantBench (Yoran et al., 2024), Mind2Web (Deng et al., 2023), WebCanvas (Pan et al., ), RealWebAssist (Ye et al., 2025), and WebLINX (Lù et al., 2024) aim to test broader agent capabilities in cross-task generalization, long-horizon reasoning, and interactive assistance. To support multimodal grounding, benchmarks like VisualWebArena (Koh et al., 2024), VideoWebArena (Jang et al., 2024), and MMINA (Zhang et al., 2024c) incorporate image and video inputs, while VisualAgentBench (Liu et al., 2024c) and VisualWebBench (Liu et al., 2024b) assess agents’ ability to align language with UI elements. However, none of these settings focus on the specialized challenges of economic data retrieval and domain-specific reasoning. While they could in principle be adapted to economics, such modifications would not ensure grounding in authoritative sources or the numeric precision required in financial contexts, which EconWebArena is designed to support. Unlike prior work that emphasizes scale or multimodal grounding, our benchmark targets domain-specific workflows that require careful handling of reporting periods and reliable numeric accuracy.
@@ 3.3 Observation and Action Spaces
EconWebArena extends the BrowserGym (Chezelles et al., 2024) framework to support economic tasks in realistic web environments. Each task provides the agent with structured observations, including the accessibility tree (AXTree), webpage screenshots, and contextual metadata such as focused elements, error logs, and action history. These inputs enable both language-based and multimodal reasoning. The action space includes fine-grained browser control primitives at both the element and coordinate levels, supporting operations like mouse movement, text input, tab switching, and form submission. This setup allows agents to perform the low-level interactions required for complex economic data retrieval. Further details are provided in Section 4.1.
@@ 4.2.3 RQ3: What directions can improve LLM agent performance on these tasks?
Our error analysis suggests several directions to improve LLM agents on EconWebArena tasks. First, robust visual grounding and multimodal reasoning are essential for interpreting economic charts, tables, and diagrams, which frequently appear in official data portals. Structured input representations, such as combining AXTree with screenshots and coordinates, can better support perception and interaction. Planning and self-criticism prompting also improve performance in multi-step navigation tasks. In addition, hybrid methods that integrate GUI-based exploration with structured API calls, when available, can enhance both reliability and efficiency. Moreover, access issues such as blocked or unstable sites remain a challenge for LLM agents, though approaches like caching snapshots or API fallbacks can help reduce their impact. Finally, domain-adaptive tuning and retrieval-augmented prompting using realistic economic examples may further strengthen agent accuracy in complex cases.
@@ B.1 More Experimental Settings
In this section, we provide additional configuration details for all experiments conducted on EconWebArena. Table 7 summarizes the prompt-level settings used by agents, adapted from the AgentLab (Chezelles et al., 2024) interface. Unless otherwise specified, these settings were fixed across all model evaluations. We evaluate a comprehensive set of LLMs, including GPT-4o555gpt-4o-2024-08-06: https://platform.openai.com/docs/models/gpt-4o (Achiam et al., 2023; OpenAI, 2024b), GPT-4o mini666gpt-4o-mini-2024-07-18: https://platform.openai.com/docs/models/gpt-4o-mini (OpenAI, 2024a), GPT-4.1777gpt-4.1-2025-04-14: https://platform.openai.com/docs/models/gpt-4.1 (OpenAI, 2025a), GPT-4.1 mini888gpt-4.1-mini-2025-04-14: https://platform.openai.com/docs/models/gpt-4.1-mini (OpenAI, 2025a), o4-mini999o4-mini-2025-04-16: https://platform.openai.com/docs/models/o4-mini (OpenAI, 2025b), Claude Sonnet 4101010claude-sonnet-4-20250514: https://openrouter.ai/anthropic/claude-sonnet-4 (Anthropic, 2025b), Gemini 2.5 Flash111111gemini-2.5-flash-preview-05-20: https://openrouter.ai/google/gemini-2.5-flash-preview-05-20 (Gemini Team, 2023; Google DeepMind, 2025), and Llama 4 Maverick121212llama-4-maverick-17b-128e-instruct-fp8: https://openrouter.ai/meta-llama/llama-4-maverick (Grattafiori et al., 2024; Meta AI, 2025).